# SuperCarver: Texture-Consistent 3D Geometry Super-Resolution for High-Fidelity Surface Detail Generation

### v1

@@ 5 Conclusion
This paper presents a specialized geometry super-resolution framework for generating high-fidelity 3D surface details as depicted in the texture appearances of given coarse meshes, serving as a powerful tool for automated surface sculpting. Our framework is normal-centric and consists of two stages. The first stage predicts view-specific detail-boosted normal maps through a deterministic prior-guided diffusion architecture. The second stage transfers the underlying geometric patterns on multi-view normal map predictions onto the original mesh surface through a noise-resistant inverse rendering scheme driven by a carefully designed distance field deformation mechanism. Experiments demonstrate the superior performance of our SuperCarver across various shape types and texture styles. We believe that our exploration can make a difference to significantly accelerate the creation of high-poly meshes, and inspire more focused research in 3D geometry super-resolution, which is highly valuable yet basically ignored in the current community.
Limitations & Future Works. Our current framework still struggles to handle elongated objects (e.g., knives, swords) because the rendered areas are too small to provide adequate appearance information under finite image resolutions. One straightforward solution is to apply localized sculpting procedures. Besides, the inherent properties of signed distance field representations restrict our capabilities to handle open surfaces and complicated topologies. Hence, we might need to introduce differentiable iso-surface extractors built upon more generic implicit geometry representations.

### v2

@@ V Conclusion
This paper presents a specialized geometry super-resolution framework for high-fidelity texture-consistent 3D surface detail generation. The proposed SuperCarver pipeline is normal-centric and composed of two sequential stages. The first stage generates detail-boosted 2D normal maps from multiple views via deterministic prior-guided normal diffusion. The second stage transfers the underlying geometric patterns of multi-view normal map predictions onto the original 3D mesh surface via noise-resistant inverse rendering driven by carefully designed distance field deformation. Extensive experiments demonstrate the effectiveness and applicability of SuperCarver across various asset types and texture styles.
Our approach serves as a powerful tool for high-poly surface sculpting and historical low-quality mesh geometry upgrading. We believe that our investigations would inspire more focused researches in the field of 3D geometry super-resolution, which is highly valuable yet rarely explored in the current community of 3D generative modeling.